OpenStudy (anonymous):

If a line contains the point (0, -1) and has a slope of 2, then which of the following points also lies on the line? A. (2, 1) B. (1, 1) C. (0, 1)

OpenStudy (anonymous):

use the point slope formula first

OpenStudy (nincompoop):

use the point-slope form

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (nincompoop):

\[y-y_1=m(x-x_1)\] where m is your slope x1 and y1 are your point coordinate

OpenStudy (nincompoop):

then solve for y

OpenStudy (anonymous):

y-(-1)=2(x-(0))

OpenStudy (anonymous):

y=2x-1

OpenStudy (nincompoop):

then do the same with the rest and whichever looks like your answer would be the answer in the choice

OpenStudy (anonymous):

graph that equation and then see which point is on that line